Calculate the area of the rectangle when a = 4 and b = 2. Next, calculate the area when the length and width are doubled.
Does the area double? What is the ratio of the two areas you calculated?
1
Expert's answer
2013-01-28T11:50:39-0500
1) The area of the rectangle is:
S1 = a*b = 4*2 = 8
2)When the length and width are doubled:
S2 = 2a*2b = 8*4 = 32
S2/S1 = 4.
So, the area is not doubled. It is multiplied by 2^2 = 4.